OpenStudy (anonymous):

(12 x -8)^2 x -8 - 9(-13+11)^2 +0 explain.

OpenStudy (studygurl14):

Can you retype it better? The way you typed it is kind of confusing and quite possibly it has missing elements. Take full advantage of parentheses. Thanks. \(\large\ddot\smile\)

OpenStudy (anonymous):

thats the best i can type it :d

OpenStudy (anonymous):

it times -8 minus 9 if thats what was confusing O.o

OpenStudy (studygurl14):

Like this?: \(\large (12\times-8)^2(-8)-9(-13+11)^2+0\)

OpenStudy (anonymous):

yeah like that

OpenStudy (studygurl14):

Okay. Do the stuff in the parenthesis (blue) first \(\large(\color{blue}{12\times-8})^2(-8)-9(\color{blue}{-13+11})^2\) (ignore the 0 on the end. It adds nothing to the expression)

OpenStudy (studygurl14):

What is \(12\times-8\), and what is \(-13+11\)?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

-96 and -2

OpenStudy (studygurl14):

Correct. So you get this: \(\large \color{green}{(-96)^2}(-8)-9\color{green}{(-2)^2}\) Next, do the stuff in green.

OpenStudy (anonymous):

so 786^2?

OpenStudy (studygurl14):

No...Remember \((-96)^2\) is another way of saying \(-96\times -96\)

OpenStudy (anonymous):

oh, well i thought i was suppose to solve it as it looks?

OpenStudy (anonymous):

OH so 9216????

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

right. and what is \((-2)^2\)

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

so then what next?

OpenStudy (anonymous):

@kyley_bear

OpenStudy (studygurl14):

no. it's not -4.

OpenStudy (anonymous):

oh right 4 lol

OpenStudy (studygurl14):

since you are multiplying -2 by itself, you have two negatives, which make a postive

OpenStudy (studygurl14):

right.

OpenStudy (studygurl14):

Now you have this: \(\large 9216(-8)-9(4)\) Do the stuff in orange \(\large(\color{orange}{9216\times -8})+(\color{orange}{-9\times 4})\)

OpenStudy (anonymous):

yeah

OpenStudy (studygurl14):

...? Did you get the answer?

OpenStudy (anonymous):

oh i zoned out lol hold on

OpenStudy (anonymous):

-73764????

OpenStudy (studygurl14):

correct

OpenStudy (anonymous):

thats a big answer though

OpenStudy (studygurl14):

lol i know

OpenStudy (anonymous):

are you sure its right?

OpenStudy (studygurl14):

yep

OpenStudy (anonymous):

oh i see